HAPTER 8 Receivables, Bad Debt Expense, and Interest Revenue LO 8-3 ?8-10 Using the Interest Formula to Compute Interest Complete the following table by computing the missing amounts (2) for the f dent cases. MB-16 Principal Amount on Note Receivable Time Period 6 months 9 months 12 months Annual Interest Rate Com off m bala 10% a. $100,000 b. $50,000 10% M8-11 Recording Note Receivable Transactions Nova Corporation hired a new product manager and agreed to provide her a $20.000 loan on a six-month, 7 percent note. Prepare journal entries to for Nova Corporation. Rather than use letters to reference each transaction, use the da LO 8-3 record the following tras transaction. The company loans the money on January 1, 2013 b. a. The new employee pays Nova the full principal and interest on its maturity date 8-3 M8-12 Recording Note Receivable Transactions Room Equipment Company received an $8,000, six-month, 6 percent note to settle a
Answer #1

Solution


a. Calculation of Interest amt of 6month, ROI 10% and principal amt is $100,000

6month Earned Interest = $100,000X10%X6/12 =$5000

Or

6 month ROI is 10%X6/12 = 5 %

6month Earned Interest = $100,000X5% = $5000

b. Calculation of Rate of Interest

Principal at $50,000

Interest Earned in 9 month period = $3,432

Convert 9 month Interest to annually = $3,432X12/9 =$4,576

Annual Interest Rate = $4,576/50,000X100 =9.152%

c. Calculation of Principal Amount

ROI 10% Annually

Earned Interest in 12 month = $4861

Principal Amount = Annual interest X100/ROI

                           = $4861X100/10 =$48,610

                                                       Table with Missing Details/Item

      Principal              ROI                   Time Period               Earned Interest

a.    $100,000                  10%                  6 Month                      $5000

b.                 $50,000                   9.152%                   9 Month                      $3,432

c.                 $48,610                   10%                   12 months                  $4,861